OpenStudy (anonymous):

given f(x) and g(x) find the indicated function f(x)=-8x+1 g(x)=4x^2+5x find f(g(-3))

OpenStudy (zzr0ck3r):

what is g(-3) ?

OpenStudy (zzr0ck3r):

I understand the question. I am asking you a question.

OpenStudy (anonymous):

sorry -3 replaces x

OpenStudy (zzr0ck3r):

\(g(x) = 4x^2+5x\), so \(g(-3) = 4(-3)^2+5(-3)=4*9-15=21\)

OpenStudy (zzr0ck3r):

so \(f(g(-3)) = f(21) = -8*21+1 = -167\)

OpenStudy (anonymous):

wouldnt you plug g into f -8(4x^2+5x-6)

OpenStudy (anonymous):

+1

OpenStudy (anonymous):

then plug in -3 for x

OpenStudy (anonymous):

is -119 right?